# Efficient Adversarial Training in LLMs with Continuous Attacks

### v1

@@ Abstract
Large language models (LLMs) are vulnerable to adversarial attacks that can bypass their safety guardrails. In many domains, adversarial training has proven to be one of the most promising methods to reliably improve robustness against such attacks. Yet, in the context of LLMs, current methods for adversarial training are hindered by the high computational costs required to perform discrete adversarial attacks at each training iteration. We address this problem by instead calculating adversarial attacks in the continuous embedding space of the LLM, which is orders of magnitudes more efficient. We propose a fast adversarial training algorithm (C-AdvUL) composed of two losses: the first makes the model robust on continuous embedding attacks computed on an adversarial behaviour dataset; the second ensures the usefulness of the final model by fine-tuning on utility data. Moreover, we introduce C-AdvIPO, an adversarial variant of IPO that does not require utility data for adversarially robust alignment. Our empirical evaluation on four models from different families (Gemma, Phi3, Mistral, Zephyr) and at different scales (2B, 3.8B, 7B) shows that both algorithms substantially enhance LLM robustness against discrete attacks (GCG, AutoDAN, PAIR), while maintaining utility. Our results demonstrate that robustness to continuous perturbations can extrapolate to discrete threat models. Thereby, we present a path toward scalable adversarial training algorithms for robustly aligning LLMs.
@@ 1 Introduction
We positively answer this research question using two novel adversarial training algorithms. We propose C-AdvUL, an efficient continuous AT algorithm, combining training on an adversarial behaviour dataset with fine-tuning on utility data. We further introduce C-AdvIPO, an adversarial variant of IPO that does not require utility data for adversarial alignment. We surpass the robustness-utility trade-offs of the discrete R2D2 AT algorithm [6], achieving up to $100\%$ attack robustness while requiring over $299$ times less computing resources. Additionally, we identify a failure mode in previous evaluation protocols: the models are tested with their chat template for safety evaluations but without it for utility evaluations. This protocol is unrealistic as the chat template is not enabled or disabled based on the prompt the user enters. By enabling the chat template for standard queries, we demonstrate that R2D2 overfits the safety objective and grammar of the harmful dataset. Thus, it often refuses to respond to benign inputs, thereby hurting its usefulness. In contrast, models trained with C-AdvUL and C-AdvIPO show substantially fewer refusals.
@@ 3 Method
In this section, we introduce our adversarial training (AT) algorithms: Continuous-Adversarial UL (C-AdvUL) and Continuous-Adversarial IPO (C-AdvIPO). We begin by reviewing the standard AT regime from Madry et al. [4] (§ 3.1). We then explain differences between attacks in the standard AT setting and unique aspects of adversarial attacks in LLMs (§ 3.2). From there, we derive the Unlikelihood loss for—C-AdvUL (§ 3.3). Next, we introduce an adversarial IPO formulation—C-AdvIPO (§ 3.5). Finally, we discuss key design decisions in the above AT algorithm (§ 3.6).
@@ 3.2 Attack Perturbation Sets in LLMs
$\delta^{t+1}=\delta^{t}+\alpha\cdot\mathrm{sign}(\nabla\log f(y|f(x+\delta^{t}))).$ (2)
@@ 3.3 Adversarial Training in LLMs
3.4 Continuous-Adversarial Unlikelihood
@@
3.5 Continuous-Adversarial IPO
Equation 3 has a similar form to DPO [30], which maximises the likelihood of a preferred answer while decreasing the likelihood of a dispreferred answer, given a prompt $x$. This motivates us to present the following loss function, which we will call Continuous-Adversarial IPO (C-AdvIPO):
@@ 3.5 Continuous-Adversarial IPO
where $\ell(h)$ would be the $\log\sigma(h)$ in the original DPO, but we use the loss proposed in Azar et al. [31] called IPO, i.e. $\ell_{\beta}(h)=\left(h-\frac{1}{2\beta}\right)^{2}$, because it is less prone to overfitting. This loss implicitly minimises the Kullback-Leibler divergence w.r.t. the original model distribution $f_{\theta_{0}}(y|x)$, which prevents the model to collapse to degenerate behaviors leading to refuse all prompts with the refusal answer $y$. As a result, we are able to omit the utility dataset for C-AdvIPO.
@@ 4 Experimental Details
For all AT experiments, we utilise the AT dataset from HarmBench [6] with the safe answer $y$ always being Sorry, I can’t do that. As a utility dataset for C-AdvUL, we employ UltraChat200k [32, 33], which has been successfully used in both the discrete AT algorithm Zephyr + R2D2 [6] and general fine-tuning [34]. For robustness evaluations, we use the first 40 samples of the HarmBench test set. Due to the substantial computational cost associated with LLM adversarial attacks, such as GCG [1], we limit our evaluation to these samples instead of the full test set.
@@
We investigate two novel continuous AT algorithms in this work C-AdvUL and C-AdvIPO. Due to the computational complexity of fine-tuning LLMs, we do not perform full model fine-tuning for both methods but use LoRA [41] on all linear layers of the transformer architectures. Additionally, we use $4$-bit quantization for all training runs to further reduce the memory overhead. We use $\ell_{2}$ norm perturbations and set the size of the attack $\epsilon$ relatively to the average magnitude of the token embeddings of the respective model. For all models, we use $10$ attack iterations. We set $\epsilon=0.1$ for Gemma and Phi-3-Mini. For Mistral-7B and Zephyr-7B, we set $\epsilon=0.05$ and $\epsilon=0.075$, respectively. For a full list of AT hyperparameters, see App. A.1.
@@ 5 Results
In Table 1, we compare the combined number of forward and backward passes used by the discrete AT algorithm RD2D [6] with C-AdvUL and C-AdvIPO. Computing a single adversarial example with R2D2 is $\approx 128.5$ times more expensive than for C-AdvUL and C-AdvIPO, while the whole training is $299$ times more costly. This illustrates the considerable compute advantage of continuous AT approaches compared to discrete methods.
@@
We first explore robustness extrapolation from continuous AT to discrete attacks for the C-AdvUL algorithm, which utilises additional utility data to maintain model performance. Figure 2 summarises the evaluation results. For all models, C-AdvUL considerably increases the average robustness against discrete adversarial attacks. For the Gemma and Zephyr models, robustness increases for all attacks. For Phi-3-Mini and Mistral-7B, PAIR still achieves high attack success rates (ASR). In terms of utility, we observe similar degradations for all C-AdvUL trained models. The MMLU and Arc scores decrease marginally, while the MT-Bench score decreases by approximately one. All models still show considerable utility after fine-tuning.
Compared to the Zephyr + R2D2 model, which was trained with discrete AT, C-AdvUL exhibits marginally worse utility on standard utility benchmarks while providing substantially improved robustness against discrete attacks. For, Zephyr + R2D2, PAIR achieves an ASR of $40\%$, while it achieves $10\%$ ASR for C-AdvUL. We note a substantial difference in the Harmless benchmark, where C-AdvUL massively outperforms Zephyr + R2D2 showing that our method has not overfitted the safety objective or the patterns in the Harmbench behaviours. Note that the Harmless score of R2D2 demonstrates that it can not simultaneously achieve non-trivial utility and robustness, which are heavily dependent on not using or using the chat template, respectively.
@@
Figure 2: Trade-off between utility and robustness for C-AdvUL (Eq. 4), C-AdvIPO (Eq. 5), and R2D2 [6], compared to their non-adversarially fine-tuned models. The objective is a small loss in utility and a large improvement in attack robustness. Larger is better for MMLU, Arc-E, Arc-C, MT-Bench (left of dashed line). Smaller is better for GCG, AutoDAN, and PAIR (right of dashed line). MT-Bench score is multiplied by 10 to see the change in performance on this $y$-axis. C-AdvIPO is not provided for 7B models in (c, d), due to computational constraints.
@@
We further investigate if adversarial variations of proven alignment methods, such as IPO, can be used to align models in an adversarially robust manner (see Figure 2). For this purpose, we fine-tune Gemma and Phi-3-Mini using the proposed C-AdvIPO algorithm. Figure 2, illustrates differences between the base model, C-AdvUL, and C-AdvIPO. Despite using no utility dataset within C-AdvIPO to retain helpfulness, the algorithm does not introduce larger utility decreases on common benchmarks than C-AdvUL. Moreover, C-AdvIPO achieves considerably higher robustness against the jailbreaking method PAIR, demonstrating generalisation to diverse threat models. The Phi-3-Mini-IPO model achieves $100\%$ attack robustness for all conducted attacks. For Gemma, robustness improvements also mostly surpass C-AdvUL, with slightly lower robustness against GCG. Compared to R2D2, C-AdvIPO does not require an auxiliary dataset to maintain utility and achieves higher robustness on average. Specifically for PAIR C-AdvIPO trained models exhibit considerably higher robustness. Lastly, the Phi-3-Mini-IPO achieves a substantially higher score on the Harmless benchmark than C-AdvUL and R2D2.
@@ 7 Adversarial Training Ablations
We found that continuous adversarial training successfully increases the robustness of LLMs to discrete adversarial attacks. Here, we explore whether robustness gains stem from using continuous adversarial attacks during training, or from the fine-tuning process itself. Thus, we fine-tune Gemma using the C-AdvIPO algorithm but without using adversarial attacks. We observe no robustness gains when fine-tuning without attacks (see App. B.2). This demonstrates that continuous adversarial attacks are a crucial part of our fine-tuning algorithm.
@@
For all our experiments, we use $10$ adversarial attack iterations. While this is orders of magnitude cheaper than calculating discrete adversarial attacks (GCG requires $2570$ model evaluations with default settings), it still increases training time by an order of magnitude. We thus propose one-step AT with C-AdvIPO. As in previous work [3], we set the step size of the attack to the magnitude of the $\epsilon$-ball. This achieves robustness improvements comparable to the multi-step variant and slightly worse utility trade-offs (see App B.1).
@@
Figure 3: Ablating how changing $\beta$ or $\epsilon{}$ affect GCG loss vs MMLU score on Gemma-IPO
@@
Prior work on AT has shown theoretical and empirical trade-offs between robustness and utility [4, 44]. Our previous results demonstrate that continuous AT can achieve non-trivial robustness-utility trade-offs. All experiments are conducted on Gemma models trained with C-AdvIPO and varying hyperparameters. Specifically, we sample $\epsilon\in[0.00125,0.3]$, and $\beta\in[0,0.5]$ and fine-tune $7$ different models. In Figure 4(b), we depict the GCG loss of the trained models (as a proxy for robustness) on the $y$-axis in logarithmic scale against the MMLU score on the $x$-axis (as a proxy for utility). Clear trade-offs between robustness and utility can be observed, ranging from models with high robustness and no utility to models showing less robustness than the standard non-robust models and slightly higher utility.
Moreover, we analyse hyperparameter choices that affect the robustness-utility trade-off for C-AdvIPO in more detail. This includes the strength of the adversarial attacks defined by the $\epsilon$ magnitude and the IPO $\beta$ value. Figure 3 illustrates that for both hyperparameters, we obtain intuitive robustness-utility trade-offs, where larger epsilon values and smaller $\beta$ values are associated with increased robustness and reduced utility. A detailed analysis can be found in App C.
@@
Figure 4: Gemma-IPO used for both plots: (a) Correlation between GCG loss and continuous attack loss. (b) GCG loss vs MMLU score for a variety of $\epsilon$ and $\beta$ values.
@@ 8 Conclusion
We answer our research question about the extrapolation of robustness under the continuous attack threat model to robustness under discrete attacks in the affirmative. We propose an efficient continuous adversarial training algorithm (C-AdvUL), combining training on an adversarial behaviour dataset with fine-tuning on utility data. Additionally, we introduce an adversarial variant of IPO (C-AdvIPO) that does not require additional utility data. Our algorithms achieve up to $100\%$ robustness against a set of state-of-the-art attacks (Phi-3-Mini-IPO), surpassing robustness utility trade-offs in previous work [6] while requiring at least $299$ times less compute. In future work, we will further analyse settings where continuous robustness does not extrapolate (e.g. novel attacks) and possible ways to address this, such as larger and more diverse training data.
@@
Our method relies on the quality and breadth of the harmful dataset, while we are less prone to overfit than Zephyr + R2D2, we may still see improvements from augmented adversarial training datasets [28]. An additional limitation is the number of hyperparameters introduced that require careful selection. We expect future work to achieve considerably better robustness-utility trade-offs through better hyperparameter selection alone. Furthermore, our proposed method C-AdvUL requires a utility dataset to retain helpfulness, which may shift the predictions of the model on unrelated tasks, a limitation we try to address with the C-AdvIPO method. Finally, due to limited compute we were not able to apply our method to much larger LLMs in the 70B parameter and larger regime, we leave this to future work.
@@ Appendix A Hyperparameter choices
Hyperparameter Gemma-UL Gemma-IPO Phi-3-Mini-UL Phi-3-Mini-IPO Mistral-7B-UL Zephyr-7B-UL

### v2

@@ Abstract
Large language models (LLMs) are vulnerable to adversarial attacks that can bypass their safety guardrails. In many domains, adversarial training has proven to be one of the most promising methods to reliably improve robustness against such attacks. Yet, in the context of LLMs, current methods for adversarial training are hindered by the high computational costs required to perform discrete adversarial attacks at each training iteration. We address this problem by instead calculating adversarial attacks in the continuous embedding space of the LLM, which is orders of magnitudes more efficient. We propose a fast adversarial training algorithm (CAT) composed of two losses: the first makes the model robust on continuous embedding attacks computed on an adversarial behaviour dataset; the second ensures the usefulness of the final model by fine-tuning on utility data. Moreover, we introduce CAPO, an adversarial variant of IPO that does not require utility data for adversarially robust alignment. Our empirical evaluation on four models from different families (Gemma, Phi3, Mistral, Zephyr) and at different scales (2B, 3.8B, 7B) shows that both algorithms substantially enhance LLM robustness against discrete attacks (GCG, AutoDAN, PAIR), while maintaining utility. Our results demonstrate that robustness to continuous perturbations can extrapolate to discrete threat models. Thereby, we present a path toward scalable adversarial training algorithms for robustly aligning LLMs.
@@ 1 Introduction
We positively answer this research question using two novel adversarial training algorithms. We propose CAT, an efficient continuous AT algorithm, combining training on an adversarial behaviour dataset with fine-tuning on utility data. We further introduce CAPO, an adversarial variant of IPO that does not require utility data for adversarial alignment. We surpass the robustness-utility trade-offs of the discrete R2D2 AT algorithm [6], achieving up to $100\%$ attack robustness while requiring over $299$ times less computing resources. Additionally, we identify a failure mode in previous evaluation protocols: the models are tested with their chat template for safety evaluations but without it for utility evaluations. This protocol is unrealistic as the chat template is not enabled or disabled based on the prompt the user enters. By enabling the chat template for standard queries, we demonstrate that R2D2 overfits the safety objective and grammar of the harmful dataset. Thus, it often refuses to respond to benign inputs, thereby hurting its usefulness. In contrast, models trained with CAT and CAPO show substantially fewer refusals.
@@ 3 Method
In this section, we introduce our adversarial training (AT) algorithms: Continuous-Adversarial Training (CAT) and Continuous-Adversarial Preference Optimisation (CAPO). We begin by reviewing the standard AT regime from Madry et al. [4] (§ 3.1). We then explain differences between attacks in the standard AT setting and unique aspects of adversarial attacks in LLMs (§ 3.2). From there, we derive the Unlikelihood loss for—CAT (§ 3.3). Next, we introduce an adversarial IPO formulation—CAPO (§ 3.5). Finally, we discuss key design decisions in the above AT algorithm (§ 3.6).
@@ 3.2 Attack Perturbation Sets in LLMs
$\delta^{t+1}=\delta^{t}+\alpha\cdot\mathrm{sign}(\nabla\log f(y|x+\delta^{t})).$ (2)
@@ 3.3 Adversarial Training in LLMs
3.4 Continuous-Adversarial Training
@@
3.5 Continuous-Adversarial Preference Optimisation
Equation 3 has a similar form to DPO [30], which maximises the likelihood of a preferred answer while decreasing the likelihood of a dispreferred answer, given a prompt $x$. This motivates us to present the following loss function, which we will call Continuous-Adversarial Preference Optimisation (CAPO):
@@ 3.5 Continuous-Adversarial Preference Optimisation
where $\ell(h)$ would be the $\log\sigma(h)$ in the original DPO, but we use the loss proposed in Azar et al. [31] called IPO, i.e. $\ell_{\beta}(h)=\left(h-\frac{1}{2\beta}\right)^{2}$, because it is less prone to overfitting. This loss implicitly minimises the Kullback-Leibler divergence w.r.t. the original model distribution $f_{\theta_{0}}(y|x)$, which prevents the model to collapse to degenerate behaviors leading to refuse all prompts with the refusal answer $y$. As a result, we are able to omit the utility dataset for CAPO.
@@ 4 Experimental Details
For all AT experiments, we utilise the AT dataset from HarmBench [6] with the safe answer $y$ always being Sorry, I can’t do that. As a utility dataset for CAT, we employ UltraChat200k [32, 33], which has been successfully used in both the discrete AT algorithm Zephyr + R2D2 [6] and general fine-tuning [34]. For robustness evaluations, we use the first 40 samples of the HarmBench test set. Due to the substantial computational cost associated with LLM adversarial attacks, such as GCG [1], we limit our evaluation to these samples instead of the full test set.
@@
We investigate two novel continuous AT algorithms in this work CAT and CAPO. Due to the computational complexity of fine-tuning LLMs, we do not perform full model fine-tuning for both methods but use LoRA [41] on all linear layers of the transformer architectures. Additionally, we use $4$-bit quantization for all training runs to further reduce the memory overhead. We use $\ell_{2}$ norm perturbations and set the size of the attack $\epsilon$ relatively to the average magnitude of the token embeddings of the respective model. For all models, we use $10$ attack iterations. We set $\epsilon=0.1$ for Gemma and Phi-3-Mini. For Mistral-7B and Zephyr-7B, we set $\epsilon=0.05$ and $\epsilon=0.075$, respectively. For a full list of AT hyperparameters, see App. A.1.
@@ 5 Results
In Table 1, we compare the combined number of forward and backward passes used by the discrete AT algorithm RD2D [6] with CAT and CAPO. Computing a single adversarial example with R2D2 is $\approx 128.5$ times more expensive than for CAT and CAPO, while the whole training is $299$ times more costly. This illustrates the considerable compute advantage of continuous AT approaches compared to discrete methods.
@@
We first explore robustness extrapolation from continuous AT to discrete attacks for the CAT algorithm, which utilises additional utility data to maintain model performance. Figure 2 summarises the evaluation results. For all models, CAT considerably increases the average robustness against discrete adversarial attacks. For the Gemma and Zephyr models, robustness increases for all attacks. For Phi-3-Mini and Mistral-7B, PAIR still achieves high attack success rates (ASR). In terms of utility, we observe similar degradations for all CAT trained models. The MMLU and Arc scores decrease marginally, while the MT-Bench score decreases by approximately one. All models still show considerable utility after fine-tuning.
Compared to the Zephyr + R2D2 model, which was trained with discrete AT, CAT exhibits marginally worse utility on standard utility benchmarks while providing substantially improved robustness against discrete attacks. For, Zephyr + R2D2, PAIR achieves an ASR of $40\%$, while it achieves $10\%$ ASR for CAT. We note a substantial difference in the Harmless benchmark, where CAT massively outperforms Zephyr + R2D2 showing that our method has not overfitted the safety objective or the patterns in the Harmbench behaviours. Note that the Harmless score of R2D2 demonstrates that it can not simultaneously achieve non-trivial utility and robustness, which are heavily dependent on not using or using the chat template, respectively.
@@
Figure 2: Trade-off between utility and robustness for CAT (Eq. 4), CAPO (Eq. 5), and R2D2 [6], compared to their non-adversarially fine-tuned models. The objective is a small loss in utility and a large improvement in attack robustness. Larger is better for MMLU, Arc-E, Arc-C, MT-Bench (left of dashed line). Smaller is better for GCG, AutoDAN, and PAIR (right of dashed line). MT-Bench score is multiplied by 10 to see the change in performance on this $y$-axis. Error bars denote $95\%$ confidence intervals, illustrating the variation in ASR estimates based on the sample size of the datasets. CAPO is not provided for 7B models in (c, d), due to computational constraints.
@@
We further investigate if adversarial variations of proven alignment methods, such as IPO, can be used to align models in an adversarially robust manner (see Figure 2). For this purpose, we fine-tune Gemma and Phi-3-Mini using the proposed CAPO algorithm. Figure 2, illustrates differences between the base model, CAT, and CAPO. Despite using no utility dataset within CAPO to retain helpfulness, the algorithm does not introduce larger utility decreases on common benchmarks than CAT. Moreover, CAPO achieves considerably higher robustness against the jailbreaking method PAIR, demonstrating generalisation to diverse threat models. The Phi-3-Mini-CAPO model achieves $100\%$ attack robustness for all conducted attacks. For Gemma, robustness improvements also mostly surpass CAT, with slightly lower robustness against GCG. Compared to R2D2, CAPO does not require an auxiliary dataset to maintain utility and achieves higher robustness on average. Specifically for PAIR CAPO trained models exhibit considerably higher robustness. Lastly, the Phi-3-Mini-CAPO achieves a substantially higher score on the Harmless benchmark than CAT and R2D2.
@@ 7 Adversarial Training Ablations
We found that continuous adversarial training successfully increases the robustness of LLMs to discrete adversarial attacks. Here, we explore whether robustness gains stem from using continuous adversarial attacks during training, or from the fine-tuning process itself. Thus, we fine-tune Gemma using the CAPO algorithm but without using adversarial attacks. We observe no robustness gains when fine-tuning without attacks (see App. B.2). This demonstrates that continuous adversarial attacks are a crucial part of our fine-tuning algorithm.
@@
For all our experiments, we use $10$ adversarial attack iterations. While this is orders of magnitude cheaper than calculating discrete adversarial attacks (GCG requires $2570$ model evaluations with default settings), it still increases training time by an order of magnitude. We thus propose one-step AT with CAPO. As in previous work [3], we set the step size of the attack to the magnitude of the $\epsilon$-ball. This achieves robustness improvements comparable to the multi-step variant and slightly worse utility trade-offs (see App B.1).
@@
Figure 3: Ablating how changing $\beta$ or $\epsilon{}$ affect GCG loss vs MMLU score on Gemma-CAPO
@@
Prior work on AT has shown theoretical and empirical trade-offs between robustness and utility [4, 44]. Our previous results demonstrate that continuous AT can achieve non-trivial robustness-utility trade-offs. All experiments are conducted on Gemma models trained with CAPO and varying hyperparameters. Specifically, we sample $\epsilon\in[0.00125,0.3]$, and $\beta\in[0,0.5]$ and fine-tune $7$ different models. In Figure 4(b), we depict the GCG loss of the trained models (as a proxy for robustness) on the $y$-axis in logarithmic scale against the MMLU score on the $x$-axis (as a proxy for utility). Clear trade-offs between robustness and utility can be observed, ranging from models with high robustness and no utility to models showing less robustness than the standard non-robust models and slightly higher utility.
Moreover, we analyse hyperparameter choices that affect the robustness-utility trade-off for CAPO in more detail. This includes the strength of the adversarial attacks defined by the $\epsilon$ magnitude and the IPO $\beta$ value. Figure 3 illustrates that for both hyperparameters, we obtain intuitive robustness-utility trade-offs, where larger epsilon values and smaller $\beta$ values are associated with increased robustness and reduced utility. A detailed analysis can be found in App C.
@@
Figure 4: Gemma-CAPO used for both plots: (a) Correlation between GCG loss and continuous attack loss. (b) GCG loss vs MMLU score for a variety of $\epsilon$ and $\beta$ values.
@@ 8 Conclusion
We answer our research question about the extrapolation of robustness under the continuous attack threat model to robustness under discrete attacks in the affirmative. We propose an efficient continuous adversarial training algorithm (CAT), combining training on an adversarial behaviour dataset with fine-tuning on utility data. Additionally, we introduce an adversarial variant of IPO (CAPO) that does not require additional utility data. Our algorithms achieve up to $100\%$ robustness against a set of state-of-the-art attacks (Phi-3-Mini-CAPO), surpassing robustness utility trade-offs in previous work [6] while requiring at least $299$ times less compute. In future work, we will further analyse settings where continuous robustness does not extrapolate (e.g. novel attacks) and possible ways to address this, such as larger and more diverse training data.
@@
Our method relies on the quality and breadth of the harmful dataset, while we are less prone to overfit than Zephyr + R2D2, we may still see improvements from augmented adversarial training datasets [28]. An additional limitation is the number of hyperparameters introduced that require careful selection. We expect future work to achieve considerably better robustness-utility trade-offs through better hyperparameter selection alone. Furthermore, our proposed method CAT requires a utility dataset to retain helpfulness, which may shift the predictions of the model on unrelated tasks, a limitation we try to address with the CAPO method. Finally, due to limited compute we were not able to apply our method to much larger LLMs in the 70B parameter and larger regime, we leave this to future work.
@@ Appendix A Hyperparameter choices
Hyperparameter Gemma-CAT Gemma-CAPO Phi-3-Mini-CAT Phi-3-Mini-CAPO Mistral-7B-CAT Zephyr-7B-CAT